OpenStudy (anonymous):

given cos sigma = -12/13, sigma in quadrant III, how would you go about to find the remaining trig functions of sigma?

OpenStudy (anonymous):

the right angled triangle with two sides of 12 and 13 is special its one which has 3 sides of lengths which are whiole numbers hypotenuse = 13 and two legs are 5 and 12

OpenStudy (anonymous):

adj side= 12 hypt = 13 so oposite side = 5 this is the triangle formed by your angle

OpenStudy (anonymous):

ok and would the sigma be the 90 degree angle then?

OpenStudy (anonymous):

sine sigma = -5/13 tan sigma = -5 / -12 = 5/12

OpenStudy (anonymous):